# Agnostic Sharpness-Aware Minimization

### v2

@@ 5 Ablation Study
5.1 Experiments on various setting of AgnosticSAM
Split training and validation set. Theoretically, the training and validation sets are independent and identically distributed samples from the data distribution. However, due to the limited number of available samples, splitting the original training set into two non-overlapping subsets for training $S^{t}$ and validation $S^{v}$ notably reduces the number of training samples. This reduction can make it challenging to train a deep neural network, particularly on small-size datasets like FGVC-Aircraft, and Flower102. To mitigate this, in practice, we randomly sample a mini-batch $B\sim S$ and then divide it into $B^{t}$ and $B^{v}$ with a ratio $|B^{t}|=r|B|$ and $|B^{v}|=(1-r)|B|$ in each iteration. As a result, the validation samples in $B^{v}$ may have been training samples in a previous iteration and the model is updated using the full training set. This approach significantly enhances the performance of Agnostic-SAM. Throughout all experiments, we use a fixed value of $r=0.7$ and use the same hyperparameters as mentioned in Section 4.1.
However, if we strictly follow the theoretical framework, where we split the original training set $S$ into two fixed training sets $S^{t}$ and validation set $S^{v}$ with $|S^{t}|=r|S|$ and $|S^{v}|=(1-r)|S|$, the algorithm still outperforms the baselines, even with considerably fewer training data. To demonstrate this, we conduct experiments under two settings: Agnostic-SAM (Non-overlap), which split $S=S^{t}\cup S^{v}$, and proposed Agnostic-SAM, which split $B=B^{t}\cup B^{v}$. Both settings use the same split ratio $r=0.7$.
In Figure 1, we present a comparative analysis of SAM and the two variants of Agnostic-SAM on CIFAR-100 (on the left) and FGVC-Aircraft (on the right). Agnostic-SAM shows notable improvement over its Non-overlap version, especially on the small-size FGVC-Aircraft dataset. However, even the Non-overlap version of Agnostic-SAM consistently outperforms SAM in most experiments.
Figure 1: Comparison of setting Agnostic-SAM (Non-overlap), proposed Agnostic-SAM, and the baseline SAM on CIFAR-100 (on the left) and FGVC-Aircraft (on the right)
Table 5: Experiments of AgnosticSAM and its simple version on CIFAR dataset.
Dataset Method WideResnet28x10 Pyramid101 Densenet121
CIFAR-10 SAM 96.72 $\pm$ 0.007 96.20 $\pm$ 0.134 91.16 $\pm$ 0.240
Agnostic-SAM (Simpler) 96.93 $\pm$ 0.021 96.26 $\pm$ 0.101 91.46 $\pm$ 0.154
Agnostic-SAM 97.17 $\pm$ 0.014 96.30 $\pm$ 0.070 91.17 $\pm$ 0.148
CIFAR-100 SAM 82.69 $\pm$ 0.035 81.26 $\pm$ 0.636 68.09 $\pm$ 0.403
Agnostic-SAM (Simpler) 83.36 $\pm$ 0.065 82.27 $\pm$ 0.311 68.97 $\pm$ 0.212
Agnostic-SAM 83.86 $\pm$ 0.077 82.93 $\pm$ 0.247 69.65 $\pm$ 0.226
Effectiveness of sharpness on validation set. As presented in Section 3.3, we first find the perturbed model $\tilde{\theta}_{l}^{v}$ by maximizing the loss on validation samples $B^{v}$, formula (4), then inject the gradient $\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ when find the actual perturbed model $\tilde{\theta}_{l}^{t}$, formula (5). In this way, the actual perturbed model $\tilde{\theta}_{l}^{t}$ maximizes the loss on training samples $B^{t}$ and minimizes the sharpness on $B^{v}$ at the same time. While the proposed method is very effective in enhancing the robustness and generalization error by minimizing sharpness on both training and validation samples, it takes extra training time per epoch.
In this section, we provide a simpler version, in which, the injected gradient when finding the actual perturbed model $\tilde{\theta}_{l}^{t}$ is $\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\theta_{l}\right)$. So that, the simpler version does not require extra steps to perturb the model on validation samples $B^{v}$ and the formula (5) becomes:
@@ 5.1 Experiments on various setting of AgnosticSAM
$\tilde{\theta}_{l}^{t}=\theta_{l}+\eta_{1}\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)-\eta_{2}\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\theta_{l}\right)$
In Table 5, we present experiments with a simpler version of Agnostic-SAM on the CIFAR dataset using various architectures. Impressively, this version consistently outperforms the baseline SAM across all settings by significant margins and achieves results comparable to those of the proposed Agnostic-SAM approach. This demonstrates the effectiveness of our method, which uses the gradient of validation samples as a guide to identify a perturbed model.
5.2 Computation complexity
The training times for SAM and Agnostic-SAM are presented in Table 6. All variations of Agnostic-SAM approach require additional computational steps to calculate the gradients on the validation set in each iteration, compared to SAM. However, in the Non-overlap version, we divide the original training set $S$ into a training set $S^{t}$ and a validation set $S^{v}$ with ratio $|S^{t}|=0.7|S|$ and $|S^{v}|=0.3|S|$, the number of training samples for this version is significantly fewer than the other methods. As a result, the training time per epoch for SAM is slightly longer than for Agnostic-SAM (Non-overlap), but considerably shorter than the other variations of Agnostic-SAM. For different versions of Agnostic-SAM, as the number of gradient steps per epoch increases, both the training time and performance tend to rise accordingly. Considering this is a trade-off for computational complexity and performance in our method.
Table 6: Compare training time (in minutes) per epoch between SAM and Agnostic-SAM on CIFAR-100 and FGVC Aircraft dataset.
Method WideResnet28x10 Pyramid101 Densenet121
SAM 2.90 $\pm$ 0.01 1.75 $\pm$ 0.00 1.40 $\pm$ 0.04
Agnostic-SAM (Non-overlap) 2.86 $\pm$ 0.00 1.72 $\pm$ 0.01 1.25 $\pm$ 0.02
Agnostic-SAM (Simpler) 3.56 $\pm$ 0.04 2.05 $\pm$ 0.01 1.90 $\pm$ 0.01
Agnostic-SAM 4.78 $\pm$ 0.02 2.86 $\pm$ 0.00 2.25 $\pm$ 0.02
EfficientNet-B4 EfficientNet-B3 ResNet18
SAM 1.41 $\pm$ 0.02 1.18 $\pm$ 0.01 0.72 $\pm$ 0.03
Agnostic-SAM (Non-overlap) 1.42 $\pm$ 0.03 1.15 $\pm$ 0.00 0.72 $\pm$ 0.01
Agnostic-SAM (Simpler) 1.75 $\pm$ 0.01 1.38 $\pm$ 0.00 0.96 $\pm$ 0.01
Agnostic-SAM 2.36 $\pm$ 0.02 1.71 $\pm$ 0.01 1.13 $\pm$ 0.00
5.3 Analysis of loss landscape
We illustrate the loss landscape of Agnostic-SAM for various architectures on the Flower102 dataset and compare it with the SAM and SGD methods. When training from scratch on EfficientNet-B4, the performance of SAM is unstable across most datasets, often requiring adjustments to several values of $\rho$ to achieve results comparable to those of SGD. This instability is reflected in the loss landscape shown in Figure 2 with high sharpness. In experiments on ResNet18, SAM updates models into wider local minima than SGD (Figure 3). In contrast, Agnostic-SAM consistently leads to a much flatter loss landscape for both the EfficientNet-B4 and ResNet18 models, indicating enhanced stability and potential for better generalization.
Figure 2: Comparing loss landscape of EffecientNet-B4 trained with (left) SGD, (middle) SAM, and (right) Agnostic-SAM on Flower102 dataset. Evidently, Agnostic-SAM leads the model to a flatter and lower loss area
Figure 3: Comparing loss landscape of ResNet18 trained with (left) SGD, (middle) SAM, and (right) Agnostic-SAM on Flower102 dataset. Both SAM and AgnosticSAM update models to flatter regions but AgnosticSAM notably achieves a smoother loss landscape compared to SAM
To further confirm that AgnosticSAM does, in fact, find minima having low curvature, we compute the Hessian for a ResNet18 trained on the Flower102 dataset with loss landscape depicted in Figure 3. The maximum eigenvalue on the log scale is 3.03 with the SGD optimizer, 2.89 with the SAM optimizer, and 2.36 with Agnostic-SAM. This is the same for the experiment with CIFAR-10 on ResNet32. Which, the maximum eigenvalue on the log scale is 4.36 with the SGD optimizer, 4.15 with the SAM optimizer, and notably lower at 3.9 with Agnostic-SAM. This provides additional evidence that Agnostic-SAM consistently leads to the discovery of minima with substantially flatter curvature compared to both SGD and SAM.
6 Conslusion
@@
In this paper, we explore the relationship between SAM and the principles underlying the MAML algorithm in terms of impact on model generalization. Inspired by MAML, we initially approach the problem of learning the best model over a training set from an agnostic viewpoint. From this perspective, we integrate sharpness-aware minimization to construct an agnostic optimization framework and introduce the Agnostic-SAM approach. This novel method optimizes the model towards wider local minima using training data, while concurrently maintaining low loss values on validation data. As a result, Agnostic-SAM is more robust against data shift issues. Through comprehensive experiments, we empirically demonstrate the superiority of Agnostic-SAM over SAM with substantial improvements in model performance across various datasets and under challenging tasks, such as when dealing with noisy labels, or various training settings on small-size datasets.
@@ Appendix A Appendix / supplemental material
In this appendix, we present the proofs in our paper.
@@ A.1 All Proofs
For any distribution $P=\mathcal{N}(\mathbf{0},\sigma_{P}^{2}\mathbb{I}_{k})$ and $Q=\mathcal{N}(\theta,\sigma^{2}\mathbb{I}_{k})$ over $\theta\in\mathbb{R}^{k}$, where $P$ is the prior distribution and $Q$ is the posterior distribution, use the PAC-Bayes theorem in [4], for all $\beta>0$, with a probability at least $1-\delta$, we have
@@
when $\|\theta\|^{2}>\sigma^{2}\big{\{}\exp(4N/k)-1\big{\}}$. Hence, for any $\|\theta\|_{2}>\sigma^{2}\big{\{}\exp(4N/k)-1\big{\}}$, we have the RHS is greater than the LHS, the inequality is trivial. In this work, we only consider the case:
@@
Thus the KL term could be bounded as follows

### v3

@@ 5 Ablation study
5.1 Cosine similarity of gradients
(a) Before updating model
(b) After updating model
(c) Improvement after updating
Figure 1: Cosine similarity of two gradients $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)$ and $\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ (a) before updating model $cosine_{b}$, (b) after updating model $cosine_{a}$ and (c) the improvement of this score $change$
In Theorem 2, we prove that minimizing the loss function $\mathcal{L}_{B^{t}}$ could encourage two gradients $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\tilde{\theta}_{l}^{t}\right)$ and $\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ to be more congruent since our update aims to maximize its lower bound, which is $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\cdot\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$. In this subsection, we measure the cosine similarity between two gradients $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)$ and $\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ before (denoted as $cosine_{b}$) and after (denoted as $cosine_{a}$) updating the model and measure the change of these two score (denoted as $change$).
@@ 5.1 Cosine similarity of gradients
$cosine_{b}$ $=\frac{\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\cdot\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\|_{2}\|\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)\|_{2}}$
$cosine_{a}$ $=\frac{\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l+1}\right)\cdot\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l+1}^{v}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l+1}\right)\|_{2}\|\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l+1}^{v}\right)\|_{2}}$
$change$ $=\frac{cosine_{a}-cosine_{b}}{cosine_{a}}$
As shown in Figure 0(c), both SAM and Agnostic-SAM improve the similarity after updating the model, this improvement also increases across training epochs. However, the similarity score of our Agnostic-SAM is always higher than SAM across the training process both before and after updating the model. This is evident that our Agnostic-SAM encourages gradient in training and validation set to be more similar during the training process.
5.2 Effectiveness of hyper-parameters
Momentum factor $\beta$.
As mentioned in section 3.3, we use momentum with a factor $\beta$ to estimate the gradients of the validation set. This approach helps stabilize the training process and ensures the model minimizes the loss across the entire validation set, rather than just a mini-batch. In this subsection, we examine the effect of the momentum factor on the model’s performance. When setting $\beta=0$, the perturbed model in each iteration maximizes the loss on a mini-batch of the training set while minimizing the loss on a mini-batch of the validation set. When $\beta>0$, the perturbed model aims to minimize the loss over the entire validation set, while maximizing the loss on a mini-batch of the training set.
The experiments are set up with the same hyper-parameters as those of experiments on CIFAR100 under noisy labels settings in Section 4.3 with basic data augmentation but without the cutout technique. We set $\rho=0.1$ for SAM and $\rho_{1}=2\rho_{2}=0.2$ for Agnostic-SAM. Results in Table 8 show that the value of $\beta$ does not significantly affect model performance overall. As such, we simply set $\beta=0.9$ in all experiments. With $\beta=0$, our method still outperforms baselines consistently, strengthening our idea of using validation gradient to indicate the model into wider local minima while reducing the generalization gap of training and testing datasets.
Table 8: Effectiveness of momentum factor $\beta$ on performance
Method SAM Agnostic-SAM
0.0 0.3 0.5 0.7 0.9
Accuracy 70.31 $\pm$ 0.2 71.14 $\pm$ 0.3 71.12 $\pm$ 0.1 70.865 $\pm$ 0.2 70.76 $\pm$ 0.3 70.91 $\pm$ 0.3
Validation batch size $|B^{v}|$ and complexity; sensitivity of perturbation radius $\rho_{1}$ and $\rho_{2}$.
Detail of these experiment is presented in Appendix A.2
6 Conclusion and Limitation
@@
In this paper, we explore the relationship between Sharpness-Aware Minimization (SAM) and the underlying principles of the Model-Agnostic Meta-Learning (MAML) algorithm, specifically in terms of their effects on model generalization. Building on this connection, we integrate sharpness-aware minimization with the agnostic perspective from MAML to develop a novel optimization framework, introducing the Agnostic-SAM approach. This method optimizes the model toward wider local minima using training data while ensuring low loss values on validation data. As a result, Agnostic-SAM demonstrates enhanced robustness against data shift issues. Through extensive experiments, we empirically show that Agnostic-SAM consistently outperforms baseline methods, delivering significant improvements in model performance across various datasets and challenging tasks. One limitation to note is that using an additional validation set when finding the perturbed model could potentially increase training time (depending on the size of the validation set). We consider this a trade-off between performance and training complexity. However, this issue could potentially be mitigated by reusing gradients from the training set in previous steps and we leave this as a direction for future work to reduce training complexity and still maintain performance.
Reproducibility Statement
We provide details of hyper-parameters for each experiment in Section 4 and Appendix A.2. Additionally, we open-source our code and provide instructions, scripts, and log files to reproduce experiments at https://anonymous.4open.science/r/AgnosticSAM-F17F/README.md
@@ Appendix A Appendix / supplemental material
In this appendix, we present the proofs in our paper and additional experiments. We open-source our code and provide instruction, scripts, and log files to reproduce experiments at https://anonymous.4open.science/r/AgnosticSAM-F17F/README.md
@@ A.1 All Proofs
For any distribution $P=\mathcal{N}(\mathbf{0},\sigma_{P}^{2}\mathbb{I}_{k})$ and $Q=\mathcal{N}(\theta,\sigma^{2}\mathbb{I}_{k})$ over $\theta\in\mathbb{R}^{k}$, where $P$ is the prior distribution and $Q$ is the posterior distribution, use the PAC-Bayes theorem in Alquier et al. (2016a), for all $\beta>0$, with a probability at least $1-\delta$, we have
@@
when $\|\theta\|^{2}>\sigma^{2}\big{\{}\exp(4N/k)-1\big{\}}$. Hence, for any $\|\theta\|_{2}>\sigma^{2}\big{\{}\exp(4N/k)-1\big{\}}$, we have the RHS is greater than the LHS, and the inequality is trivial. In this work, we only consider the case:
@@
Thus the KL term could be bounded as follow
@@
A.2 Additional Experiments
Validation batch size $|B^{v}|$ and complexity
Our method is to use a gradient on the validation set as a helper indicator to lead the model to wider local minima while maintaining low loss on the validation set, and the model should be updated mainly using training samples. Increasing validation mini-batch size could potentially increase performance and training time. In Table 9, we present the results of Agnostic-SAM with various validation batch sizes $|B^{v}|$ of CIFAR-100 with Resnet32 while maintaining a fixed training batch size $|B^{t}|=512$, the other hyper-parameters are the same as above experiments with momentum factor $\beta$. We consider performance and training complexity to be the trade-off of Agnostic-SAM and find that setting $|B^{t}|=4|B^{v}|$ works well for all experiments.
Table 9: Experiments on different sizes of validation mini-batch with a fixed size of training mini-batch is 512 samples
Method Validation Accuracy Training time
batch-size (s/epochs)
SAM 0 70.31 $\pm$ 0.233 11s
Agnostic-SAM 16 70.58 $\pm$ 0.219 11s
32 71.07 $\pm$ 0.212 12s
64 70.67 $\pm$ 0.049 13s
128 71.21 $\pm$ 0.056 14s
256 71.04 $\pm$ 0.219 15s
Sensitivity of perturbation radius $\rho_{1}$ and $\rho_{2}$
Throughout this paper, we used a consistent setting of $\rho_{1}=2\rho_{2}=2\rho$, where $\rho$ represents the perturbation radius in the SAM method for all experiments. While these hyperparameters could be optimized for each experiment individually, we find that this configuration delivers good performance across most experiments. By setting $\rho_{1}>\rho_{2}$, we ensure that the perturbed model prioritizes maximizing the loss on the training set rather than minimizing it on the validation set. This approach encourages the model to focus primarily on minimizing sharpness during the actual update step in Formula 6.
To verify the impact of these hyperparameters on model performance, we conduct experiments with varying perturbation radius and present the results in Figure 2. Notably, the configuration where $\rho_{1}>\rho_{2}$ consistently yields higher accuracy compared to the setting where $\rho_{1}<\rho_{2}$. When increasing $\rho_{2}$, the model places more emphasis on minimizing the validation set loss, rather than sharpness on the training set during the actual update step in Formula 6. This shift in focus can lead to overfitting, ultimately reducing performance.
Figure 2: Experiments of various perturbation radius $\rho_{1}$ and $\rho_{2}$
Analysis of loss landscape and eigenvalues of the Hessian matrix
We demonstrate the effectiveness of Agnostic-SAM in guiding models toward flatter regions of the loss landscape, as compared to both SAM and SGD, in Figures 3 and 4. The loss landscapes are visualized with the same setting, the blue areas represent lower loss values, while the red areas indicate higher loss values. Although SAM is shown to lead the model to a flatter region than SGD, Agnostic-SAM achieves an even smoother and significantly flatter loss landscape, especially in experiments with EfficientNet-B2 in Figure 3.
Figure 3: Loss landscape of EffecientNet-B2 trained on Flower102 dataset with (left) SGD, (middle) SAM, and (right) Agnostic-SAM.
Figure 4: Loss landscape of ResNet32 trained (left) SGD, (middle) SAM, and (right) Agnostic-SAM on Cifar100 dataset.
Methods Top-5 eigenvalues of Hessian matrix
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $\lambda_{5}$
EfficientNet-B2 on Flower102
SGD $2.05\times 10^{5}$ $0.45\times 10^{5}$ $0.26\times 10^{5}$ $-0.47\times 10^{5}$ $-0.49\times 10^{5}$
SAM $1.61\times 10^{3}$ $1.34\times 10^{3}$ $1.23\times 10^{3}$ $1.04\times 10^{3}$ $-0.97\times 10^{3}$
Agnostic-SAM $0.61\times 10^{3}$ $0.41\times 10^{3}$ $0.37\times 10^{3}$ $0.32\times 10^{3}$ $0.31\times 10^{3}$
Resnet32 on Cifar100
SGD $3.07\times 10^{6}$ $2.40\times 10^{6}$ $2.10\times 10^{6}$ $1.64\times 10^{6}$ $1.46\times 10^{6}$
SAM $1.50\times 10^{6}$ $1.14\times 10^{6}$ $0.96\times 10^{6}$ $0.87\times 10^{6}$ $0.81\times 10^{6}$
Agnostic-SAM $1.04\times 10^{6}$ $0.79\times 10^{6}$ $0.66\times 10^{6}$ $0.58\times 10^{6}$ $0.57\times 10^{6}$
Table 10: Eigenvalues of Hessian matrix
To further validate that Agnostic-SAM successfully locates minima with low curvature, we compute the Hessian of the loss landscape and report the five largest eigenvalues, sorted from $\lambda_{1}$ to $\lambda_{5}$, in Table 10. These eigenvalues provide insight into the curvature of the model at the optimized parameters. Larger eigenvalues indicate steeper curvature, meaning the model is more sensitive to small changes in its parameters. Conversely, smaller eigenvalues suggest flatter minima, which are typically associated with improved robustness, better generalization, and reduced sensitivity to overfitting. Negative eigenvalues indicate non-convex curvature in certain directions.
As shown in Table 10, Agnostic-SAM consistently achieves positive and lower eigenvalues compared to the baseline methods, suggesting that it effectively leads the model toward flatter regions of the loss landscape. These results further support the efficacy of Agnostic-SAM in optimizing for smoother and more stable solutions across a variety of architectures and tasks.